OpenStudy (aakashsudhakar):

(Gravitational Field Problem) A spacecraft makes a trip from the Earth to the Moon, 380,000km (3.8E8m) away. At what point in the trip will the gravitational field be zero? The mass of the Moon is about (1/81) that of the Earth.

OpenStudy (sid123):

motion is influenced by gravity or motion can exists independetly

OpenStudy (anonymous):

Things not influenced by gravity have motion, ie, photons; gravity is just an energy sink around which things agglomerate?

OpenStudy (sid123):

can u describe the term gravity is just energy sink around which things agglomerate?

OpenStudy (anonymous):

Erf mün systm  The gravitational field should be zero where the force of the Earth's gravitational field exactly equals that of the Moon's. \[ R = 3.8 x 10^8m \\ \ \\ \ r = \textrm{distance from moon to zero grav point}\] \[ \sum F_r = -F_m + F_e = 0 \\ \ \\ \frac{G\frac{m}{81}m_o}{r^2} = \frac{Gmm_o}{(R-r)^2} \\ \ \\ \ \\ \frac{r^2}{(R-r)^2} = \frac{1}{81} \\ \ \\ \ \] Solve for r Look up Lagrange point #1 for a much more refined answer and approach ^_^